• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ramesh Subedi

    Log entry time 16:02:39 on August11,2006

    Entry number 179162

    keyword=Shift summary

    08:00 Shift starts with run number 2521 (left arem) leftover from the owl shift.

    Target Liquid Hydrogen
    Spectrometer angles: 23.902 (left), 28.283 (right)
    Kinematics 15: elastic production on LH2 at p = 0.5211 GeV/c.



    Run Start Run (left / right) Events (left)

    08:07 2522 / 20881 1.9 M
    08:30 2523 / 20882 1.9 M
    08:56 2524 / 20883 1.9 M
    09:23 2525 / 20884 1.9 M
    09:46 2526 / 20885 1.9 M

    10:09 Kinematics 16: elastic production on LH2 at p = 0.5107 GeV/c.
    Left HRS Magnet IOC rebooted from the web (http://hareboot10) since Q1 and Q3 do not accept the new momentum setting.

    Runs 2527 and 2528 (left) are junk while setting prescale factors.

    PS3 changes from 2 to 3, other prescale factors remain as before.

    10:43 2529 / 20888 1.9 M
    10:55 2530 / 20889 1.9 M
    11:06 2531 / 20890 1.9 M
    11:21 2532 / 20891 1.9 M
    11:34 2533 / 20892 1.9 M
    11:47 2534 / 20893 1.9 M
    11:59 2535 / 20894 1.9 M
    12:12 2536 / 20895 1.9 M
    12:24 2537 / 20896 1.9 M
    12:39 2538 / 20897 1.9 M
    12:52 2539 / 20898 1.9 M

    13: 04 Kinematics 17: elastic production on LH2 at p = 0.5007 GeV/c.

    13:08 2540 / 20899 1.9 M
    13:18 2541 / 20900 1.9 M
    13:30 2542 / 20901 1.9 M
    13:41 2543 / 20902 1.9 M
    13:52 2544 / 20903 1.9 M
    14:05 2545 / 20904 1.9 M
    14:16 2546 / 20905 junk (left hrs coda crashed)
    14:32 2547 / 20906 1.9 M

    14:54 2549 / 20908 junk (beam stopped)
    14:55 controlled excess for radiation and other EH&S stuff check in the hall

    15:41 beam permit

    15:51 beam back
    15:51 2550 / 20909 1.9 M